According to Newton’s law of cooling, the rate at which a substance cools in moving
air is proportional to the difference between the temperature of the substance and that
of the air. If the temperature of the air is C
o
290 and the substance cools from
C
o
370 to C
o
330 in 10 minutes, find when the temperature will be C
o
295 .
1
Expert's answer
2020-09-14T11:16:01-0400

By the condition "\frac{dT}{dt}\backsim(T-T_a)"

where "\frac{dT}{dt}" - the rate of cooling; "T"- the temperature of substance; "T_a" - the temperature of air.

We can rewrite this expression as "\frac{dT}{dt}=-k(T-T_a)" , where "k" - coefficient of proportionality. Solve this equation.

"\frac{dT}{dt}=-k(T-T_a)"

"\frac{dT}{T-T_a}=-kdt"

"\int_{T_1}^{T_2}\frac{dT}{T-T_a}=-k\int_{0}^{t} dt"

where "T_1=370", "T_2=330", "t=10", "T_a=290".

"ln\frac{T_1-T_a}{T_2-T_a}=kt"

"ln\frac{370-290}{330-290}=ln\frac{80}{40}=ln2"

"ln2=10k"

"k=0.07"

Now, find the low of cooling of temperature.

"\int\frac{dT}{T-T_a}=-0.07\int dt"

"ln(T-T_a)=-0.07t"

"e^{ln(T-T_a)}=e^{-0.07t}"

"T-T_a=e^{-0.07t}"

"T=T_a+e^{-0.07t}"

Substitute "T_3=295".

"295=290+e^{-0.07t}"

"e^{-0.07t}=5"

"-0.07t=ln5"

"t=23"

Answer: t=23 minutes.